Cost of Trade Credit

A large retailer obtains merchandise under the credit terms of 2/15, net 35, but routinely takes 55 days to pay its bills. (Because the retailer is an important customer, suppliers allow the firm to stretch its credit terms.) What is the retailer's effective cost of trade credit? Assume 365 days in year for your calculations. Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

Calculate the effective cost of trade credit as follows:

Effective cost of credit = (1+ (Discount rate /(1 - Discount rate))^n - 1

Effective cost of credit = (1+(2%/(1-2%))^(365/(55-15)) - 1

Effective cost of credit =( 102.0408163265306%)^(365/40)) - 1

Effective cost of credit = 1.2024 -1

Effective cost of credit = 0.2024 or 20.24%.
Answer #2

Cost of Trade Credit

A large retailer obtains merchandise under the credit terms of 2/15, net 40, but routinely takes 65 days to pay its bills. (Because the retailer is an important customer, suppliers allow the firm to stretch its credit terms.) What is the retailer's effective cost of trade credit? Assume a 365-day year. Do not round intermediate calculations. Round your answer to two decimal places.


source: g
answered by: f